OpenStudy (anonymous):

-3 square root of 24 * 6 square root of 12 (simplify)

OpenStudy (anonymous):

\[-3\sqrt{24}\times6\sqrt{12}\]\[-18\sqrt{12\times12\times2}\]\[-18\times12\sqrt{2}\]\[-216\sqrt{2}\]